answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

True or false p lies in plane b the line containing p and q must lie in plane b?

False.

In order for the line PQ to lie in plane B, then both P and Q must lie in plane B.

If two numbers are complementary what is their probability?

On the basis that numbers are a continuous variable, the probability of any particular number (or pair) is 0.

How many outcomes can you get if you make seven number random phone numbers?

The answer depends on what you are measuring. The random variable could be:

  • whether or not the phone rings (27 = 128 outcomes)
  • whether or not the phone is answered (128 outcomes)
  • whether a male or female answers (37 = 2187 outcomes - including no answer).

Other variables will generate other outcomes.

What is the probability of a 5 not being rolled on a dice?

Probability is desired options over total options.

There are 6 faces on a standard dice, so NOT rolling a 5 is 5/6.

When three dice are thrown together what is the probability of getting a total score of 10?

As there are 3 dice, it is quite easy (and quick) to work out the possible ways a sum of 10 can be made:

If the first die is a 1, the second and third dice could be {3, 6}; {4, 5}; {5, 4}; {6, 3} - 4 possibilities

If the first die is a 2, the second and third dice could be {2, 6}; {3, 5}; ...; {6, 2} - 5 possibilities

If the first die is a 3, the second and third dice could be {1, 6}; {2, 5}; ...; {6, 1} - 6 possibilities

If the first die is a 4, the second and third dice could be {1, 5}; {2, 4}; ...; {5, 1} - 5 possibilities

If the first die is a 5, the second and third dice could be {1, 4}; {2, 3}; ...; {4, 1} - 4 possibilities

If the first die is a 6, the second and third dice could be {1, 3}; {2, 2}; {3, 1} - 3 possibilities

So there are 4 + 5 + 6 + 5 + 4 + 3 = 27 possible ways of getting a sum of 10

With 3 dice there are 6 × 6 × 6 = 216 possible ways of them landing

→ pr(sum 10) = 27/216 = 1/8 = 0.125 = 12.5%

How is probability used in genetics?

probability is used in genetics to determine the possibilities of offspring having a certain trait

What is the probabililty of at least 2 people same birthday from a group of 13 people?

19.4%

CALCULATION:

The probability of at least 2 people having the same birthday in a group of 13

people is equal to one minus the probability of non of the 13 people having the

same birthday.

Now, lets estimate the probability of non of the 13 people having the same birthday.

(We will not consider 'leap year' for simplicity, plus it's effect on result is minimum)

1. We select the 1st person. Good!.

2. We select the 2nd person. The probability that he doesn't share the same

birthday with the 1st person is: 364/365.

3. We select the 3rd person. The probability that he doesn't share the same

birthday with 1st and 2nd persons given that the 1st and 2nd don't share the same

birthday is: 363/365.

4. And so forth until we select the 13th person. The probability that he doesn't

share birthday with the previous 12 persons given that they also don't share

birthdays among them is: 353/365.

5. Then the probability that non of the 13 people share birthdays is:

P(non of 13 share bd) = (364/365)(363/365)(362/365)∙∙∙(354/365)(353/365)

P(non of 13 share bd) ≈ 0.805589724...

Finally, the probability that at least 2 people share a birthday in a group of 13

people is ≈ 1 - 0.80558... ≈ 0.194 ≈ 19.4%

The above expression can be generalized to give the probability of at least x =2

people sharing a birthday in a group of n people as:

P(x≥2,n) = 1 - (1/365)n [365!/(365-n)!]

What is a patent bet?

it is a full cover bet is any bet which consists of all available multiple bets over a given number of selections.

What is meant by the term probability?

Probability means - The likelihood that a particular event will occur.

What is the probability of picking a queen from a pack of playing cards?

  1. In a standard 52 card deck: 4 queens / 52 cards = about 7.69 %
  2. In a deck with the jokers: 4 queens / 54 cards = about 7.41 %
  3. In a Pinochle deck: 8 queens / 48 cards = about 16.67 %
  4. In a standard modern Tarot deck: 4 queens / 78 cards = about 5.13 %
  5. etc.

What is the probability of choosing something to wear?

depends on choices. 4 pair of shoes -- pick 1. 6 shirts pick 1. etc

1/4 x 1/6 x etc

What is the difference between 'probability' and 'proportion'?

  • well a proportion compares part of a quanity to the whole quanity called the base using a percent
  • probabiltyis the cahnce of an event happening for example like getting a puppy that is a simple event
Trending Questions
What is the probability that the sum of the numbers on two dice is even when they are rolled? 60 percent pass the 1st time and 40 percent fail 2nd time around 80 percent pass and 20 percent fail 3rd time 90 percent pass and 10 percent fail what is the probability you pass on the 2nd time? What is the face value of a jack in a deck of cards? A type of selection where the normal distribution is pushed apart into two separate peaks? Do you have sets of multiple choice questions for arnis? What probability value would be needed to complete the following probability distribution? How many 3 cards are in a deck? Tristan guesses on two multiple-choice questions on a test if each question has five possibe answers choices what is the probability that he gets the first one correct and the second one incorrect? How many different sums of money can be made using at most one bill of each of the folloing denominations 1 5 10 20 50? What is the chance of picking a red king out of a pack of cards? Can a union steward be a supervisor? Why is the sample mean an unbiased estimator of the population mean? You want working model on thermocol sheet related to maths like on the topic geometry etc? If you flip two coins what is the probability that they will both land on tails? Where can you buy jelly marbles? What 3 ways probability can be shown? What is the standard deviation and what is the probability that a student scored higher than 322 if the mean test score is 312 and 16 percent of the scores are under 307? What is the probaility of drawing a 10 from a standard deck? What is stratified sampling? 13 cards are drawn simultaneously from a pack of 52 cardsif aces count 1face cards 10 and other according to denominationfind the expectation of the total score on 13 cards?